A woman is six times as old as her son. In 9 years she will be three times as old as her son. How old are they now?
Answer by srinivas.g(540)

let woman age be x now. Her son's age be y now
woman is six times as old as her son.
x =6*y
x =6y
In 9 years she will be three times as old as her son
so, after 9 years her age will be x+9 & her son's age will be y+9
so, x+9 = 3*(y+9)
put x=6y in the above equation
6y+9= =3*(y+9)
6y+9= 3y+ 3*9
6y+9 = 3y+27
subtract 3y on both sides
6y+9 -3y= 3y+27-3y
3y+9= 27
subtract 9 on both sides
3y+9-9=27-9
3y= 18
divide with 3 on both sides
3y/3 = 18/3
y= 6
so age of son now (y) = 6
age of woman (x=6y) = 6*6 =36
age of son is 6 years and woman is 36 years